# Fast emulation of fermionic circuits with matrix product states

### v2

@@ I Introduction
For fermionic problems in particular, the Fermionic Quantum Emulator (FQE) is a software package that aims for maximum efficiency in full statevector emulation [12]. The FQE implements algorithms designed specifically to take full advantage of symmetries that are common to fermionic systems, in particular the particle number and projected spin symmetries. This source of efficiency enables emulation of quantum circuits that target fermionic simulation at larger scales than are achievable with naive statevector emulation. FQE provides near-optimal performance for most emulation tasks on single-node, CPU platforms. In recent work [12], it was shown that exploiting just number and projected spin-symmetries can lead to substantial savings in total run-time and memory requirements over qubit based simulations that do not use these symmetries.
@@ II Theoretical Background
where indexes the physical occupations of the th orbital (or lattice site). Contracting the matrix dimensions (“virtual bonds”) of all the matrices for a given vector of physical occupations, will produce the amplitude associated with that basis state. This representation is clearly exact if the “bond dimension” is allowed to be arbitrarily large. In fact, the MPS tensors can be obtained constructively by repeated singular value decomposition of the exact coefficient tensor. The dimension of the virtual bonds required for an exact representation grows exponentially toward the center of the tensor product so that the upper bound of the virtual bond summation for the th tensor is obtained as for spin-1/2 fermions. As a result, in order to benefit from the MPS Ansatz, one must place a truncated upper bound on the bond dimension, , so that . The effectiveness of the truncated MPS can be understood from information theoretic arguments [31], and it is known that the optimal approximate (truncated) wavefunction is obtained by retaining the states in a Schmidt decomposition with the largest Schmidt numbers [33]. Note that construction of an MPS by consecutive SVDs of the FCI tensor does not yield the optimal truncated MPS wavefunction. Applying a local variational optimization to the truncated MPS Ansatz results in the density matrix renormalization group (DMRG) algorithm. Physical symmetries can be efficiently treated within the MPS representation by blocking the MPS tensors by local particle number and projected spin labels and then only storing those blocks which correctly multiply to the conserved global symmetry [34].
@@
The bond dimension of the MPO is determined by the number of non-zero terms in the expansion and by the locality of the operators in each term. The efficient construction of MPOs of dense operators is a non-trivial problem, and the MPS-FQE can, via its interface with pyblock3, use either an SVD-based compression approach [35] or the “bipartite” approach [36]. As with the MPS, the MPOs can be blocked by their symmetry labels. This is vitally important for efficient computation of expectation values and renormalized basis states used in the DMRG and time-dependent DMRG (td-DMRG) algorithms [35]. Given a bra and ket MPS, the matrix element of the MPO can be exactly computed with an efficient contraction algorithm which is at the heart of the utility of DMRG. Unitary evolution of the MPS wavefunction is a more difficult problem and it has been an active research topic over the last 20 years. For the purposes of quantum emulation, we want to be able to apply the unitary operators of arbitrary, but generally sparse, generators. One option is to use the 4th order Runge-Kutta integrator (RK4),
@@ IV.1 Preparing near-exact ground states
Many quantum algorithms for the ground state of fermionic systems use some variant of quantum phase estimation (QPE) which requires a significant overlap between the initial state, , and the exact ground state, . MPS-FQE can be used both as a means to approximate the exact ground state, , and as a way to emulate state preparation algorithms for . In the former case, DMRG is the most efficient way to form near-exact states in the MPS representation, and a seamless interface with the pyblock3 and block2 codes makes it easy to convert the DMRG solution to an MPS-FQE wavefunction. For example, in Figure 2 we show the potential energy and the overlap, , for a restricted Hartree–Fock (RHF) initial state of N. The “exact” quantities are computed from extrapolating DMRG results to the limit of infinite bond dimension. In this particular example, the overlap between the RHF state and the “exact” ground state is large enough such that the RHF state could probably be used for QPE over most of the potential energy surface. Note that it is useful to be able to compute these quantities for larger basis sets, as we do here, because minimal basis sets that are often used while discussing quantum algorithms are too small to provide quantitative information with chemically-meaningful accuracy. The calculations with the 6-31G basis set are still possible using an exact statevector solver, but the def2-SVP basis is far too large for an exact solution.
MPS-FQE can also be used to emulate state preparation algorithms on quantum computers. For example, one might want to use a small amount of quantum imaginary time evolution or adiabatic state preparation to prepare a state that has stronger overlap with the exact ground state. In this case, the qualitative information provided by MPS-FQE with a modest bond dimension can be very useful. In Figure 3 we compare the amount of imaginary and real time respectively necessary to prepare a state where the overlap with the exact ground state is greater than 0.9. Again we show results for N () in the def2-SVP basis set where the 28 spatial orbitals (56 spin orbitals) makes exact emulation impossible. Although imaginary time evolution is much more efficient in terms of the propagation time, implementation on a quantum computer is more complicated and the most efficient algorithm will depend on the implementation on real hardware. Either way, MPS-FQE can provide useful estimates such applications.
@@ IV.2 Testing VQE Ansätze
The ansatz obtained via the ADAPT-VQE algorithm is, in principle, exact for a given basis and can be expressed as
@@
In Figure 4, we present ADAPT-VQE results obtained with MPS-FQE for N using the 6-31G basis set. For the calculations described here, we have evaluated the operator pool gradients using exact analytical commutation relations (although all Hamiltonian applications are performed using an approximate sweep algorithm) despite the fact that any unitary corresponding to a selected operator is ultimately applied approximately under truncated bond dimension. The ADAPT ansatz is subsequently optimized at each iteration using finite difference evaluations to remain consistent with the approximate propagation scheme. While a conventional ADAPT-VQE implementation considers convergence with respect to the norm of the operator pool gradient vector [64], we have found in this application that one cannot converge the gradient norm below about for (compared to in the ADAPT() implementation) using the described scheme for evaluating gradients with respect to parameters for operators in the pool.
@@
The reason the operator pool gradient norm remains relatively large is likely due to the mismatch between the manner in which the gradients are evaluated and the selected unitaries are applied. The operator pool gradient vector is computed assuming that the application of each corresponding unitary will be done exactly (i.e. they are obtained via the analytical commutation relations). The quality of this assumption likely deteriorates as the ADAPT ansatz grows and the impact of a truncated bond dimension becomes more severe for larger iterations. Thus, it remains possible that an operator selected from the pool, which is expected to have the greatest influence on the energy under the assumption of an exact application of the corresponding unitary, may have a considerably smaller impact under truncated bond dimension.
@@
An alternative, and perhaps more internally consistent, implementation would evaluate the operator pool gradients via finite differences, such that they reflect the approximate nature of propagation in a truncated Hilbert space. However, while we perform the ansatz optimization during each iteration in this way, evaluating the gradients of every term in the expansive operator pool via finite differences would be prohibitively expensive for the basis set considered here. The data shown in Figure 4 therefore corresponds to the endpoint of an ADAPT trajectory that is terminated when the cumulative change in energy over 3 successive iterations is less than .
@@
This emulation of ADAPT-VQE contains both error from the approximate nature of the ADAPT-VQE method and error due to the truncated bond dimension. The former includes the error due to the finite number of iterations used to the optimize the ansatz and the possibility of converging to a local minimum. Unfortunately, these two sources of error are coupled, which makes extrapolation to the “exact” ADAPT-VQE energy difficult. For example, the ADAPT-VQE solution at may be a different solution than the one found at . Despite these difficulties, MPS emulation in the fermionic representation allows one to study the qualitative features of the ansatz for much larger systems. In this case, the basis of 36 spin orbitals is much larger than systems studied with ADAPT-VQE in the past.
@@ IV.3 Numerical evaluation of Trotter error
When deciding on the most efficient strategy for real applications, an asymptotic understanding of the error is not sufficient and the coefficients must be also considered. For example, for the simple Lie–Trotter formula,

### v4

@@ I Introduction
For fermionic problems in particular, the Fermionic Quantum Emulator (FQE) is a software package that aims for maximum efficiency in full statevector emulation [12]. The FQE implements algorithms designed specifically to take full advantage of symmetries that are common to fermionic systems, in particular the particle number and projected spin symmetries. This source of efficiency enables emulation of quantum circuits that target fermionic simulation at larger scales than are achievable with naive statevector emulation. FQE provides near-optimal performance for most emulation tasks on single-node, CPU platforms. In a recent work [12], it was shown that exploiting just number and projected spin-symmetries can lead to substantial savings in total run-time and memory requirements over qubit based simulations that do not use these symmetries.
@@ II Theoretical Background
where indexes the physical occupations of the th orbital (or lattice site). Contracting the matrix dimensions (“virtual bonds”) of all the matrices for a given vector of physical occupations, will produce the amplitude associated with that basis state. This representation is clearly exact if the “bond dimension" is allowed to be arbitrarily large. In fact, the MPS tensors can be obtained constructively by repeated singular value decomposition of the exact coefficient tensor. The dimension of the virtual bonds required for an exact representation grows exponentially toward the center of the tensor product so that the upper bound of the virtual bond summation for the th tensor is obtained as for an -site system of spin-1/2 fermions. As a result, in order to benefit from the MPS Ansatz, one must place a truncated upper bound on the bond dimension, , so that . The effectiveness of the truncated MPS can be understood from information theoretic arguments [31], and it is known that the optimal approximate (truncated) wavefunction is obtained by retaining the states in a Schmidt decomposition with the largest Schmidt numbers [33]. Note that construction of an MPS by consecutive SVDs of the FCI tensor does not yield the optimal truncated MPS wavefunction. Applying a local variational optimization to the truncated MPS Ansatz results in the density matrix renormalization group (DMRG) algorithm. Physical symmetries can be efficiently treated within the MPS representation by blocking the MPS tensors by local particle number and projected spin labels and then only storing those blocks which correctly multiply to the conserved global symmetry [34].
@@
The bond dimension of the MPO is determined by the number of non-zero terms in the expansion and by the locality of the operators in each term. The efficient construction of MPOs of dense operators is a non-trivial problem, and the MPS-FQE can, via its interface with pyblock3, use either an SVD-based compression approach [35] or the “bipartite" approach [36]. As with the MPS, the MPOs can be blocked by their symmetry labels. This is vitally important for efficient computation of expectation values and renormalized basis states used in the DMRG and time-dependent DMRG (td-DMRG) algorithms [35]. Given a bra and ket MPS, the matrix element of the MPO can be exactly computed with an efficient contraction algorithm which is at the heart of the utility of DMRG. Unitary evolution of the MPS wavefunction is a more difficult problem and it has been an active research topic over the last 20 years. For the purposes of quantum emulation, we want to be able to apply the unitary operators of arbitrary, but generally sparse, generators. One option is to use the 4th order Runge-Kutta integrator (RK4),
@@ IV.1 Preparing near-exact ground states
Many quantum algorithms for the ground state of fermionic systems use some variant of quantum phase estimation (QPE) which requires a significant overlap between the initial state, , and the exact ground state, . MPS-FQE can be used both as a means to approximate the exact ground state, , and as a way to emulate state preparation algorithms for . In the former case, DMRG is the most efficient way to form near-exact states in the MPS representation, and a seamless interface with the pyblock3 and block2 codes makes it easy to convert the DMRG solution to an MPS-FQE wavefunction. For example, in Figure 2 we show the potential energy and the overlap, , for a restricted Hartree–Fock (RHF) initial state of N2. The “exact” quantities are computed from extrapolating DMRG results to the limit of infinite bond dimension. In this particular example, the overlap between the RHF state and the “exact" ground state is large enough such that the RHF state could be used for QPE over most of the potential energy surface. Note that it is useful to be able to compute these quantities for larger basis sets, as we do here, because minimal basis sets that are often used while discussing quantum algorithms are too small to provide quantitative information with chemically-meaningful accuracy. The calculations with the 6-31G basis set are still possible using an exact statevector solver, but the def2-SVP basis is far too large for an exact solution.
MPS-FQE can also be used to emulate state preparation algorithms on quantum computers. For example, one might want to use a small amount of quantum imaginary time evolution or adiabatic state preparation to prepare a state that has stronger overlap with the exact ground state. In this case, the qualitative information provided by MPS-FQE with a modest bond dimension can be very useful. In Figure 3 we compare the amount of imaginary and real time respectively necessary to prepare a state where the overlap with the exact ground state is greater than 0.9. Note that the “exact" ground state is estimated by extrapolating DMRG calculations to zero truncation error from . Again we show results for N2 () in the def2-SVP basis set where the 28 spatial orbitals (56 spin orbitals) makes exact emulation impossible. Although imaginary time evolution is much more efficient in terms of the number of time steps, each step of imaginary time is much more complicated to implement on a quantum computer [59, 60] and the most efficient algorithm will depend on the implementation on real hardware. Either way, MPS-FQE can provide useful estimates for such applications.
@@ IV.2 Testing VQE Ansätze
The ansatz obtained via the ADAPT-VQE algorithm is, in principle, exact for a given basis [66] and can be expressed to th order as
@@
In Figure 4, we present ADAPT-VQE results obtained with MPS-FQE for N2 using the 6-31G basis set with a max bond dimension of . For the calculations described here, we have evaluated the operator pool gradients using exact analytical commutation relations (although all Hamiltonian applications are performed using an approximate sweep algorithm) despite the fact that any unitary corresponding to a selected operator is ultimately applied approximately via RK4 followed by compression under the imposed maximum bond dimension and a cutoff value of . The ADAPT ansatz is subsequently optimized at each iteration using analytical commutation relations. We also note that, with respect to the exact potential energy curve, one obtains significantly better agreement simply by performing DMRG simulations, representing a direct variational optimization of the MPS ansatz. The ADAPT-VQE ansatz, on the other hand, may be much more sensitive to orbital ordering given the nature of the applied excitation operators. In fact, we have performed the same set of calculations with and observed only a very small (almost graphically indistinguishable) improvement to the energy. While we have not utilized any orbital localization or reordering schemes in this application, we note that the MPS-FQE library is well-suited for evaluating their impact in this context, and it is reasonable to expect that the use of such schemes would improve approximate MPS-based ADAPT-VQE results.
@@
While a conventional ADAPT-VQE implementation considers convergence with respect to the norm of the operator pool gradient vector [66], we have found in this application that it can be challenging to converge the gradient norm below about in many cases (compared to in the ADAPT() implementation) using the described scheme for evaluating gradients with respect to parameters for operators in the pool. The reason the operator pool gradient norm remains relatively large in some cases is likely due to the mismatch between the manner in which the gradients are evaluated and the selected unitaries are applied. The operator pool gradient vector is computed assuming that the application of each corresponding unitary will be done exactly (i.e. they are obtained via the analytical commutation relations). The quality of this assumption likely deteriorates as the ADAPT ansatz grows and the impact of a truncated bond dimension becomes more severe for larger iterations. Thus, it remains possible that an operator selected from the pool, which is expected to have the greatest influence on the energy under the assumption of an exact application of the corresponding unitary, may have a considerably smaller impact under truncated bond dimension.
@@
An alternative, and perhaps more internally consistent, implementation would evaluate all gradients via finite differences, such that they reflect the approximate nature of propagation in a truncated Hilbert space. However, evaluating the gradients of every term in the expansive operator pool via finite differences would be prohibitively expensive for the basis set considered here. Analytic evaluation of the gradient of the approximate ansatz is also possible in principle, but the implementation is difficult, requiring the solution of response equations at each step. The data shown in Figure 4 therefore corresponds to the endpoint of an ADAPT trajectory that is terminated when the cumulative change in energy over 3 successive iterations is less than or, in some cases, when the gradient norm converges below .
@@
This emulation of ADAPT-VQE contains error from the approximate nature of the ADAPT-VQE method, error due to the truncated bond dimension, and error due to the adopted convergence scheme. The former includes the error due to the finite number of iterations used to optimize the ansatz as well as the possibility of converging to a local minimum. Unfortunately, these sources of error are coupled, which makes extrapolation to the “exact" ADAPT-VQE energy difficult. For example, the ADAPT-VQE solution at may be a different solution than the one found at . Despite these difficulties, MPS emulation in the fermionic representation allows one to study the qualitative features of the ansatz for much larger systems. In this case, the basis of 36 spin orbitals is much larger than systems studied with ADAPT-VQE in the past.
@@ IV.3 Numerical evaluation of Trotter error
To decide on the most efficient strategy for real applications, an asymptotic understanding of the error is not sufficient and the coefficients must be also considered. For example, for the simple Lie–Trotter formula,